# Stealthy Poisoning Attacks Bypass Defenses in Regression Settings

### v1

@@ Abstract
Regression models are widely used in industrial processes, engineering and in natural and physical sciences, yet their robustness to poisoning has received less attention. When it has, studies often assume unrealistic threat models and are thus less useful in practice. In this paper, we propose a novel optimal stealthy attack formulation that considers different degrees of detectability and show that it bypasses state-of-the-art defenses. We further propose a new methodology based on normalization of objectives to evaluate different trade-offs between effectiveness and detectability. Finally, we develop a novel defense (BayesClean) against stealthy attacks. BayesClean improves on previous defenses when attacks are stealthy and the number of poisoning points is significant.
@@ I Introduction
Despite this, the robustness of regression models to poisoning attacks has not been widely studied. Most studies that examine it [44, 68, 35, 49, 64, 62, 38], assume the attacker’s objective is solely to maximize the damage, without considering attacks may aim to be stealthy i.e., less detectable. In this paper, we show that the defenses proposed in these studies fail to mitigate stealthy attacks.
We argue that in practical scenarios, attackers will seek to remain undetected as well as to maximize the attack effectiveness because if poisoning is detected, the training set may be discarded thwarting the attack entirely. In contrast, a stealthy attack may be more costly and less effective, but nevertheless damage to the model significantly.
@@
To answer the first question, we propose a novel formulation of poisoning attacks as a parameterizable combination of attack objectives: (i) maximizing the damage to the model and (ii) minimizing the risk of being detected. We model the attacker’s problem as a novel multiobjective bilevel optimization problem and show how it can be solved. This formulation is more general than adaptive attacks, which target specific weaknesses of the defense, and allows us to construct attacks that simultaneously evade different types of defenses. For this, we exploit attack transferability, and show that the appropriate modeling of detectability constraints, targeting a standard model, is enough to bypass the defenses. Furthermore, we show that in some cases defenses are not only unable to mitigate the attack, but also damage the model’s performance compared to an unprotected model. In other words, use of these defenses can be detrimental.
We focus on Linear Regression (LR), as most existing literature focuses on this setting. LR is also widely used in practice so more important to defend [42, 35, 48, 62]. However, the analysis of linear models also allow us to gain more general insights about attack effectiveness and limitations of defenses. We demonstrate these insights by also evaluating the performance and scalability of our attack in nonlinear models like Deep Neural Networks (DNNs).
Our results show that noise and uncertainty about predictions play a pivotal role in the success of poisoning attacks. Consequently, to answer RQ2 and overcome limitations of existing defenses, we propose BayesClean, a novel defense based on Bayesian LR. We show that BayesClean is capable of reducing the impact of stealthy poisoning attacks, by rejecting points based on their location within the predictive variance of the model. This principle is more promising for rejecting suspicious points compared to previous approaches and can form the basis for the development of further defenses.
In summary, the contributions of this paper are:
  • We propose a novel stealthy attack formulation against regression algorithms via multiobjective bilevel optimization: one objective accounts for attack effectiveness, the other for the detectability of poisoned points. For this, we introduce a novel detectability-risk function and a novel approach for solving the corresponding optimization problem.
  • We show that state-of-the-art defenses fail to mitigate the stealthy attacks we propose, even when they are not adaptive. We evaluate the different defenses proposed so far in the literature on real-world datasets, in both LR models and DNNs, and identify the factors that influence their effectiveness. To our knowledge, we are the first to analyze optimal data poisoning attacks targeting DNNs in regression settings.
@@ II Related Work
In terms of defenses, estimation procedures for robust regression that are not strongly affected by the presence of stochastic noise and outliers are available in robust statistics [34, 18]. However, these methods fail in high dimensions and struggle to identify carefully crafted poisoning samples [67]. Liu et al. [39] propose robust LR algorithms when the feature matrix has low rank and can be projected to a lower dimensional space. Diakonikolas et al. [15] propose SEVER, a robust meta-algorithm that can detect poisoning samples through outlier scores based on the singular value decomposition of the gradients. Weerasinghe et al. [62] propose BIG-LID, an LK defense for nonlinear regression learning based on the concept of Local Intrinsic Dimensionality (LID) [^1] The LID value of a data point identifies the minimum number of latent variables required to represent that data point and recent evidence suggests that LID is an indicator of the degree of being an outlier [29, 63]. [30] to identify and reduce the impact of poisoning samples. All these methods have provable robustness guarantees, but the assumptions on which they rely are not usually satisfied in practical settings where attackers also consider attack detectability.
@@
Finally, Debugging Using Trusted Items (DUTI) [68] identifies samples that have the biggest influence on the validation error on a small trusted clean set.
As a final remark, note that most of these defenses assume that they know the fraction of poisoning samples in the training dataset. This is unrealistic in practical applications. Furthermore, deviations on the estimation of the fraction of poisoning points can negatively impact the model’s performance for both poisoned and clean models. In contrast, our BayesClean defense does not require estimating the fraction of poisoning points to reject, as we reject points according to the predictive variance of a Bayesian LR model.
@@ III Threat Model
where the first term represents the MSE evaluated on the training set and $\Omega({\bf w})$ is the regularization term controlled by the hyperparameter $\lambda$ commonly used to prevent overfitting. $\lambda$ can be adjusted by using cross validation or by minimizing the MSE on a separate validation dataset. In this paper we leverage $L_{2}$ regularization, i.e., ridge regression.
@@ III-B Attacker’s Knowledge
We consider an attacker aware of the training data, the feature set, the model, and the objective function—a worst case scenario. However, we do not assume knowledge of the defense used by the victim. Thus, we are in a Limited Knowledge (LK) threat model where the attacker targets an unprotected regression model, and we evaluate the attack effectiveness against models that use different defenses. We chose this setting because we aim to use a more general formulation of the poisoning attack that takes into account the risk of being detected. In contrast, defensive algorithms proposed in the literature differ in their assumptions or assume adaptive attacks where the attacker has knowledge of the specific defense deployed. For example, DUTI, requires the defender to have access to a clean validation set, whereas other approaches, like TRIM, do not. Given the heterogeneity of these methods, rather than crafting different adaptive attacks targeting specific defenses, we use the same attack algorithm to show that such defenses fail to detect and mitigate stealthy attacks that seek to evade detection. While stealthy attacks may be sub-optimal compared to adaptive attacks, they nevertheless allow attackers to reach their goal. Our results show that our general formulation allows us to bypass existing defenses, and, shockingly, that in some cases, defenses lead to worse performance than an unprotected model.
@@ IV-A Optimal Attack Formulation with Soft Detectability Constraints
There is no general agreement on how to define the detectability risk (or detectability constraints). Hard constraints, [^3] In an optimization problem, a hard constraint is a constraint that must be satisfied by any feasible solution to the model. On the other hand, a soft constraint can be violated, but violating the constraint incurs a penalty in the objective function (often, the greater the amount by which the constraint is violated, the greater the penalty) [46]. typically considered in adversarial ML [59, 14, 33, 31, 23, 19, 37, 12], must be satisfied by any feasible solution. In practice, the attacker’s willingness to reduce attack effectiveness to avoid detection depends on the attacker’s risk appetite. In particular, the adversary may not be aware of the defense algorithm. Even if the defense algorithm is known and has a concrete threshold for removing points [35, 15], the rejection region can be very complex.
@@
The hyperparameter $\alpha$ can be set by the attacker based on its risk appetite and a parametric sweep allows us to investigate different effectiveness detectability trade-offs. Note that, when combining objectives in this way, the correlation between the objective functions is known to affect the search space structure and algorithm performance. The outer objective $\mathcal{A}_{d}$ should not be convex or concave, otherwise it can be difficult to achieve a smooth transition for this trade-off. In the extreme case where $\mathcal{R}\propto$-$\mathcal{L}$, the multiobjective problem reduces to a single-objective problem where it is not possible to find a trade-off.
@@ IV-A1 Modeling the Detectability-Risk Function
To evaluate the detectability risk, i.e., the attacker’s perception of the risk of being detected, different formulations are possible, reflecting the different a defender can detect poisoned points.
@@ IV-A3 Normalization of Objective Functions
For the effectiveness objective, the normalization coefficient depends on the size of the batch of points, and on whether the objective is measured on the clean or the poisoned model. Therefore, it is difficult to obtain the maximum value of the objective, $\max_{\mathcal{D}_{\mathrm{p}}^{\prime}\in\Phi(\mathcal{D}_{\mathrm{p}})}\mathcal{L}(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star})|_{\mathcal{R}=0}$, as this value depends on the values of the previous batches of poisoning points, which may, in turn, be constrained. To address this challenge, we calculate a new normalization coefficient for each batch i.e., for the $i$th batch, we re-normalize the effectiveness objective as follows: $\mathcal{L}_{\mathrm{norm}}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)$ $\approx{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)}/{\mathcal{L}_{\mathrm{ref}}}$ $\approx\frac{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)}{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i-1)}}\right)}$, where ${\bf w}^{\star^{(i-1)}}$ are the poisoned parameters trained on the previous optimal poisoned set. When computing the first batch, the previous poisoned set is the clean training set, and therefore ${\bf w}^{\star^{(i-1)}}={\bf w}^{\star}_{\mathrm{tr}}$. Although this heuristic does not calculate the maximum of the objective, it provides a good empirical reference for the range of values the effectiveness objective can take. Different values of $\alpha$ lead to different normalization coefficients see Alg. 2 (Appx. A).
In summary, to solve the multiobjective bilevel optimization problem (4), we first compute the normalization coefficient for the detectability-risk function for $\alpha=1$ (Alg. 1), and then, for $\alpha\neq 1$, we compute the corresponding normalization coefficient for the effectiveness function (Alg. 2). Fig. 3 shows the effect of normalizing the objective functions in a synthetic example. Note that normalization enables a smoother transition as the value of $\alpha$ changes. Fig. 4 shows that poisoning points close to the clean distribution ($\alpha=0.4$) shift the regression line even when TRIM is deployed as a defense. In contrast, poisoning points without detectability constraints ($\alpha=1$) are easily filtered by TRIM.
@@ IV-C Experimental Evaluation of Stealthy Attacks
We now empirically show that our attack formulation with soft detectability constraints bypasses state-of-the-art defenses. For this, we evaluate the effectiveness and stealth of poisoning attacks in stationary settings. Thus, we perform a random split of the datasets to simulate absence of concept drift (i.e., changes in the data distribution through time) [^5] A time-aware split, rather than a random split would make it difficult to determine whether the effectiveness of the poisoning attack is due to an weakness of the regression model or due to the natural evolution of data [56].. We use four real-world datasets: interest rate of loans (Loan)—made on the Lending Club peer-to-peer lending platform [36] (preprocessed as in [35]); heart disease (Heart Disease) [62]; Boston housing prediction (Boston Housing) [28]; and appliances energy prediction (Appliances) [8] (preprocessed as in [62]). These are also the datasets used in the related work on poisoning attacks or in standard regression benchmarks. All datasets are drawn at random from the original joint pool of points and details for each dataset are included in Table I (Appx. E).
@@
We evaluate our stealthy attack against several benchmark defenses: TRIM [35], Huber [34], and SEVER [15] testing attacks with different levels of detectability, i.e., for different values of $\alpha$. Thus, we examine the trade-off between effectiveness and stealthiness: $\alpha=1$ are attacks entirely dominated by effectiveness, $\alpha=0.3$ are stealthy attacks that achieve some effectiveness, and $\alpha=0.1$ are attacks dominated by stealthiness.
@@ V-C BayesClean
V-D Experimental Evalutation of BayesClean
@@ V-D2 Deep Neural Networks
For completeness, we also show in Fig. 11 the performance of our defense on DNNs. Even though BayesClean is based on a Bayesian LR model, its predictive variance is able to detect poisoning points targeting DNNs. As in the case of LR, our defense is competitive in most cases, being particularly useful when the attacker considers detectability and the ratio of poisoning points is large. It is also evident, according to the negative values of the test defense gain, that defending DNNs is more challenging. These results suggest further research is needed on methods to defend DNNs in regression settings.

### v2

# arXiv:2601.22308v2[cs.LG] 02 Mar 2026

@@ Abstract
Regression models are widely used in industrial processes, engineering, and in natural and physical sciences, yet their robustness to poisoning has received less attention. When it has, studies often assume unrealistic threat models and are thus less useful in practice. In this paper, we propose a novel optimal stealthy attack formulation that considers different degrees of detectability and show that it bypasses state-of-the-art defenses. We further propose a new methodology based on normalization of objectives to evaluate different trade-offs between effectiveness and detectability. Finally, we develop a novel defense (BayesClean) against stealthy attacks. BayesClean improves on previous defenses when attacks are stealthy and the number of poisoning points is significant.
@@ I Introduction
Despite this, the robustness of regression models to poisoning attacks has not been widely studied. Most studies on this topic [44, 68, 35, 49, 64, 62, 38], assume the attacker’s objective is solely to maximize the damage, without considering attacks may aim to be stealthy, i.e., less detectable. In this paper, we show that the defenses proposed in these studies fail to mitigate stealthy attacks. We argue that in practical scenarios, attackers will seek to maximize the attack effectiveness while remaining undetected, because if poisoning is detected, the training set may be discarded thwarting the attack entirely. Although stealthy attacks may be more costly and less effective compared to unconstrained attacks, they can produce a significant more damage in practice as they are more difficult to detect.
@@
To answer the first question, we propose a novel formulation of poisoning attacks as a parameterizable combination of attack objectives: (i) maximizing the damage to the model and (ii) minimizing the risk of being detected. We model the attacker’s problem as a multiobjective bilevel optimization problem and show how it can be solved. This formulation is more general than adaptive attacks, which target specific weaknesses of the defense, and allows us to construct attacks that simultaneously evade different types of defenses. For this, we exploit attack transferability, and show that the appropriate modeling of detectability constraints, targeting a standard model, is enough to bypass the defenses. Furthermore, we show that in some cases defenses are not only unable to mitigate the attack, but also damage the model’s performance compared to an unprotected model. In other words, the use of these defenses can be detrimental.
We focus on Linear Regression (LR), as most existing literature focuses on this setting. LR is also widely used in practice, so it is important to defend these models [42, 35, 48, 62]. The analysis of the robustness of linear models also allows us to gain more general insights about attack effectiveness and limitations of existing defense strategies. Beyond LR, we also demonstrate these insights by also evaluating the performance and scalability of our attack in nonlinear models like Deep Neural Networks (DNNs).
Our results show that noise and uncertainty about predictions play a pivotal role in the success of stealthy poisoning attacks. Thus, based on this observation we propose BayesClean a novel defense capable of overcoming some of the limitations of existing defenses to mitigate stealthy attacks and, consequently, answering RQ2. We show that BayesClean is capable of reducing the impact of stealthy poisoning attacks, by rejecting points based on their location within the predictive variance of the model. This principle is more promising for detecting and rejecting suspicious points compared to previous approaches and can form the basis for the development of further defenses.
In summary, the contributions of this paper are as follows:
  • We propose a novel stealthy poisoning attack formulation against regression models via multiobjective bilevel optimization: One objective accounts for attack effectiveness, the other for the detectability of poisoned points. For this, we introduce a novel detectability-risk function and a novel approach for solving the optimization problem.
  • We show that state-of-the-art defenses fail to mitigate the stealthy attacks we propose, even when they are not adaptive. We evaluate the different defenses proposed so far in the literature on real-world datasets, in both LR models and DNNs, and identify the factors that influence their effectiveness. To our knowledge, we are the first to analyze optimal indiscriminate poisoning attacks targeting DNNs in regression settings.
@@ II Related Work
In terms of defenses, algorithms for robust regression that are not strongly affected by the presence of stochastic noise and outliers are available in robust statistics [34, 18]. However, these methods fail in high dimensions and struggle to identify carefully crafted poisoning samples [67]. Liu et al. [39] propose robust LR algorithms when the feature matrix has low rank and can be projected to a lower dimensional space. Diakonikolas et al. [15] propose SEVER, a robust meta-algorithm that can detect poisoning samples through outlier scores based on the singular value decomposition of the gradients. Weerasinghe et al. [62] propose BIG-LID, an LK defense for nonlinear regression learning based on the concept of Local Intrinsic Dimensionality (LID) [^1] The LID value of a data point identifies the minimum number of latent variables required to represent that data point and recent evidence suggests that LID is an indicator of the degree of being an outlier [29, 63]. [30] to identify and reduce the impact of poisoning samples. All these methods have provable robustness guarantees, but the assumptions on which they rely are not usually satisfied in practical settings where attackers also consider attack detectability.
@@
Finally, Debugging Using Trusted Items (DUTI) [68] identifies samples that have the biggest influence on the validation error evaluated on a small trusted clean set.
It is important to note that most of these defenses assume that they know the fraction of poisoning samples in the training dataset. This is unrealistic in practical applications. Furthermore, deviations on the estimation of the fraction of poisoning points can negatively impact the model’s performance for both poisoned and clean models. In contrast, our BayesClean defense does not require estimating the fraction of poisoning points to reject, as we reject points according to the predictive variance of a Bayesian LR model.
@@ III Threat Model
where the first term represents the MSE evaluated on the training set and $\Omega({\bf w})$ is the regularization term controlled by the hyperparameter $\lambda$ commonly used to prevent overfitting. $\lambda$ can be adjusted by using cross validation or by minimizing the MSE on a separate validation dataset. In this paper we use $L_{2}$ regularization, i.e., ridge regression.
@@ III-B Attacker’s Knowledge
We consider an attacker aware of the training data, the feature set, the model, and the objective function—a worst case scenario. However, we do not assume knowledge of the defense used by the victim. Thus, we are in a Limited Knowledge (LK) threat model where the attacker targets an unprotected regression model, and we evaluate the attack effectiveness against models that use different defenses. We chose this setting because we aim to use a more general formulation of the poisoning attack that takes into account the risk of being detected. In contrast, defensive algorithms proposed in the literature differ in their assumptions or assume adaptive attacks where the attacker has knowledge of the specific defense deployed. For example, DUTI, requires the defender to have access to a clean validation set, whereas other approaches, like TRIM, do not. Given the heterogeneity of these methods, rather than crafting different adaptive attacks targeting specific defenses, we use the same attack algorithm to show that such defenses fail to detect and mitigate stealthy attacks that seek to evade detection. While stealthy attacks may be sub-optimal compared to adaptive attacks in Perfect Knowledge (PK) settings, they nevertheless allow attackers to reach their goal, especially in LK scenarios where attackers are not aware of the target model and the defense used. Our results show that our general formulation allows us to bypass existing defenses, and, shockingly, that in some cases, defenses lead to worse performance than an unprotected model.
@@ IV-A Optimal Attack Formulation with Soft Detectability Constraints
There is no general agreement on how to define the detectability risk (or detectability constraints). Hard constraints, [^3] In an optimization problem, a hard constraint is a constraint that must be satisfied by any feasible solution to the model. On the other hand, a soft constraint can be violated, but violating the constraint incurs a penalty in the objective function (often, the greater the amount by which the constraint is violated, the greater the penalty) [46]. typically considered in adversarial ML literature [59, 14, 33, 31, 23, 19, 37, 12], must be satisfied by any feasible solution. In practice, the attacker’s willingness to reduce attack effectiveness to avoid detection depends on the attacker’s risk appetite. In particular, the adversary may not be aware of the defense algorithm. Even if the defense algorithm is known and has a concrete threshold for removing points [35, 15], the rejection region can be very complex.
@@
The hyperparameter $\alpha$ can be set by the attacker based on its risk appetite and a parametric sweep allows us to investigate different effectiveness detectability trade-offs. Note that, when combining objectives in this way, the correlation between the objective functions is known to affect the search space structure and algorithm performance. To enable a smooth trade-off, the outer objective $\mathcal{A}_{d}$ should admit multiple local optima rather than a single one.
@@ IV-A1 Modeling the Detectability-Risk Function
To evaluate the detectability risk, i.e., the attacker’s perception of the risk of being detected, different formulations are possible, as defenders can aim to detect poisoning points in two different ways:
@@ IV-A3 Normalization of Objective Functions
For the effectiveness objective, the normalization coefficient depends on the size of the batch of points, and on whether the objective is measured on the clean or the poisoned model. Therefore, it is difficult to obtain the maximum value of the objective, $\max_{\mathcal{D}_{\mathrm{p}}^{\prime}\in\Phi(\mathcal{D}_{\mathrm{p}})}\mathcal{L}(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star})|_{\mathcal{R}=0}$, as this value depends on the values of the previous batches of poisoning points, which may, in turn, be constrained. To address this challenge, we calculate a new normalization coefficient for each batch i.e., for the $i$th batch, we re-normalize the effectiveness objective as follows: $\mathcal{L}_{\mathrm{norm}}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)$ $\approx{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)}/{\mathcal{L}_{\mathrm{ref}}}$ $\approx\frac{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i)}}\right)}{\mathcal{L}\left(\mathcal{D}_{\mathrm{val}},{\bf w}^{\star^{(i-1)}}\right)}$, where ${\bf w}^{\star^{(i-1)}}$ are the poisoned parameters trained on the previous optimal poisoned set. When computing the first batch, the previous poisoned set is the clean training set, and therefore ${\bf w}^{\star^{(i-1)}}={\bf w}^{\star}_{\mathrm{tr}}$. Although this heuristic does not calculate the maximum of the objective, it provides a good empirical reference for the range of values the effectiveness objective can take. Different values of $\alpha$ lead to different normalization coefficients see Alg. 2 (Appx. A). e In summary, to solve the multiobjective bilevel optimization problem (4), we first compute the normalization coefficient for the detectability-risk function for $\alpha=1$ (Alg. 1), and then, for $\alpha\neq 1$, we compute the corresponding normalization coefficient for the effectiveness function (Alg. 2). Fig. 3 shows the effect of normalizing the objective functions in a synthetic example. Note that normalization enables a smoother transition as the value of $\alpha$ changes. Fig. 4 shows that poisoning points close to the clean distribution ($\alpha=0.4$) shift the regression line even when TRIM is deployed as a defense. In contrast, poisoning points without detectability constraints ($\alpha=1$) are easily filtered by TRIM.
@@ IV-C Experimental Evaluation of Stealthy Attacks
We now empirically show that our attack formulation with soft detectability constraints bypasses state-of-the-art defenses. For this, we evaluate the effectiveness and stealth of poisoning attacks in stationary settings. Thus, we perform a random split of the datasets to simulate absence of concept drift. [^5] A time-aware split, rather than a random split would make it difficult to determine whether the effectiveness of the poisoning attack is due to an weakness of the regression model or due to the natural evolution of data [56]. We use four real-world datasets: interest rate of loans (Loan)—made on the Lending Club peer-to-peer lending platform [36] (preprocessed as in [35]); heart disease (Heart Disease) [62]; Boston housing prediction (Boston Housing) [28]; and appliances energy prediction (Appliances) [8] (preprocessed as in [62]). These are also the datasets used in the related work on poisoning attacks or in standard regression benchmarks. All datasets are drawn at random from the original joint pool of points and details for each dataset are included in Table I (Appx. E).
@@
We evaluate our attack against several benchmark defenses: TRIM [35], Huber [34], and SEVER [15] testing attacks with different levels of detectability, i.e., for different values of $\alpha$. We examine the trade-off between effectiveness and stealthiness: $\alpha=1$ are attacks entirely dominated by effectiveness, $\alpha=0.3$ are stealthy attacks that achieve some effectiveness, and $\alpha=0.1$ are attacks dominated by stealthiness.
@@ V-C BayesClean
V-D Experimental Evaluation of BayesClean
@@ V-D2 Deep Neural Networks
For completeness, in Fig. 11 we show the performance of our defense on DNNs. Even though BayesClean is based on a Bayesian LR model, its predictive variance is able to detect poisoning points targeting DNNs. As in the case of LR, our defense is competitive in most cases, being particularly useful when the attacker considers detectability and the ratio of poisoning points is large. It is also evident, according to the negative values of the test defense gain, that defending DNNs is more challenging. These results suggest further research is needed on methods to defend DNNs in regression settings.